OpenStudy (anonymous):

How does (1-x)-1 equal (x-2) Here are my issues. 1. Can I arrange the equation -1(1-x)? If not why not? I am confused by some algebra rules here. 2. How can I just switch the (1-x) to (x-1). In (1-x) the 1 is positive, and in (x-1) its negative??? 3. How do I just subtract the -1 from whats inside the parens? Don't I get (1-x-1)? Is this equal to (2-x)?

OpenStudy (anonymous):

in general it doesn't

OpenStudy (anonymous):

Can you number your answer so I know which question you are addressing? (Even though they are all related haha)

OpenStudy (anonymous):

i am assuming by ! you do not mean factorial. \(1-x-1=-x\) and \(-x=x-2\iff -2x=-2\iff x=2\)

OpenStudy (anonymous):

Yes lol let me fix that

OpenStudy (anonymous):

it is not clear to me what the question is if it is "solve for \(x\): \(1-x-1=x-2\) then the answer is \(x=2\)

OpenStudy (anonymous):

oh Im taking the derivative, but Im just simplifiying here

OpenStudy (anonymous):

Im taking the derivative of (1-x)e^-x and one of the parts I am simplifying is (1-x)-1 And I don't get exactly how and why (1-x)-1 equals (x-2)